OpenStudy (anonymous):

Linear Algebra: Let A be a 4x3 matrix and suppose that the vectors z1=(1,1,2)^T and (1,0,-1)^T form a basis for N(A). If the b = a1 + 2a2 + a3 (where b is a vector and a1, a2, a3 are column vectors of A), find all solutions of the system Ax = b.

OpenStudy (anonymous):

z2 = (1,0,-1)^T *correction

OpenStudy (jamesj):

So first of all we know that Az1 = Az2 = 0. Hence for any solution x = y of Ax = b, then x = y + c1.z1 + c2.z2 is also a solution for any scalars c1 and c2.

OpenStudy (jamesj):

Now the question is what is such a vector y? Given the form of b, you have been given a serious clue by the question. Look carefully at the form of b and figure out what the entries of the column vector y would have te be to get Ay = b = a1 + 2a2 + a3

OpenStudy (jamesj):

If you're really stuck, ask yourself this: if a1, a2 and a3 are the column of A, then what, for instance is the result of A(1,0,0)^T? And what about A(0,1,0)^T?

OpenStudy (anonymous):

i'm sorry if this sounds really dumb, but could you please explain how were you able to say Az1 = Az2 = 0 from the fact that {z1,z2} is a basis for N(A)?

OpenStudy (zarkon):

read the definition of null space

OpenStudy (anonymous):

Please tell me if this is a correct thought process: {z1, z2} is a basis for N(A). Therefore any linear combination of z1 and z2 will be z : Az = 0 Therefore if I let scalar c1=1 and c2=0, c1z1 will be z : Az=0 The same can be said for z2 if c1=0 and c2=1. Therefore Az1 = Az2 = 0

OpenStudy (zarkon):

{z1, z2} is a basis for N(A). clearly z1 and z2 are in the nullspace thus Az1=0 and Az2=0

OpenStudy (anonymous):

Lol i really have tendencies of making everything complicated. Thank you.

OpenStudy (anonymous):

Thank you all, I understand the solution now. My professor already gave the class a solution, but I couldn't make sense of how the reminiscent method of solving non-homogeneous ODE's could appear here. You all gave me another way of looking at this problem, which makes more sense to me. Thanks.